Find the amount that results from the investment.
$12,000 invested at 11% compounded quarterly after a period of 4 years

Answer by mananth(12270)

Principal P = 12000
Amount= A
years=n 4.00
compounded 4 times a year t
Rate = 11.00 0.11
Amount = P*((n+r)/n)^n*t

Amount = = 12000 *( 1 + 0.11 /t)^ 4 * 4
Amount = 12000 *( 1 + 0.03 )^ 16
12000 *( 1.03 )^ 16
Amount = $18522.11